# FlowerTune: A Cross-Domain Benchmark for Federated Fine-Tuning of Large Language Models

### v1

@@ 6 Conclusion
Limitations and future work. This study focuses primarily on cross-silo federated settings; future work may explore a broader range of scenarios, including more resource-constrained environments and more heterogeneous data distributions. Additionally, the experiments were conducted using a unified configuration following a systematic protocol but without extensive hyper-parameter tuning for each individual model-dataset pair. We envision this work as a catalyst for developing more inclusive, secure, and domain-adapted language models, and invite the broader research community to contribute to and extend this benchmarking effort.
@@ Appendix A Detailed experimental settings
This section provides details on the base models (Section A.1) and hyper-parameter configurations (Section A.2) used in the experiments.
A.1 Base model configurations
@@
Non-Instruct Mistral-7B-v0.3 [5] 7 B 33
Mistral-Small-24B-Base-2501 [3] 24 B 131
Gemma-2-9B [55] 9 B 256
Phi-4 [1] 14 B 100
Llama-3.2-1B [23] 1 B 128
Llama-3.2-3B [23] 3 B 128
Llama-3.1-8B [23] 8 B 128
Qwen2-0.5B [17] 0.5 B 152
Qwen2.5-1.5B [63] 1.5 B 152
Qwen2.5-3B [63] 3 B 152
Qwen2.5-7B [63] 7 B 152
SmolLM2-135M [7] 135 M 49
SmolLM2-360M [7] 360 M 49
SmolLM2-1.7B [7] 1.7 B 49
Instruct Mistral-7B-Instruct-v0.3 [6] 7 B 33
Mistral-Small-24B-Instruct-2501 [4] 24 B 131
Gemma-2-9B-Instruct [48] 9 B 256
Phi-4-Mini-Instruct [2] 3.8 B 200
Llama-3.2-1B-Instruct [23] 1 B 128
Llama-3.2-3B-Instruct [23] 3 B 128
Llama-3.1-8B-Instruct [23] 8 B 128
Qwen2-0.5B-Instruct [17] 0.5 B 152
Qwen2.5-1.5B-Instruct [63] 1.5 B 152
Qwen2.5-3B-Instruct [63] 3 B 152
Qwen2.5-7B-Instruct [63] 7 B 152
SmolLM2-135M-Instruct [7] 135 M 49
SmolLM2-360M-Instruct [7] 360 M 49
SmolLM2-1.7B-Instruct [7] 1.7 B 49
A.2 Hyper-parameter configurations
@@
Appendix B Additional experimental results
B.1 Federated fine-tuning on non-instruct pre-trained base models
@@
B.2 Pareto analysis of evaluation performance
Figure 4: Accuracy (%) versus system performance for different non-instruct base models federated fine-tuned on the General NLP challenge, presented in table 11. The errorbar indicates ±$\pm 1$ Std. Dev. on the different downstream tasks.
Figure 5: Accuracy (%) versus system performance for different non-instruct base models federated fine-tuned on the Finance challenge, presented in table 12. The errorbar indicates ±$\pm 1$ Std. Dev. on the different downstream tasks.
Figure 6: Accuracy (%) versus system performance for different non-instruct base models federated fine-tuned on the Medical challenge, presented in table 13. The errorbar indicates ±$\pm 1$ Std. Dev. on the different downstream tasks.
Figure 7: Pass@1 scores (%) versus system performance for different non-instruct base models federated fine-tuned on the Coding challenge, presented in table 14. The errorbar indicates ±$\pm 1$ Std. Dev. on the different downstream tasks.
Figure 8: Accuracy (%) versus system performance for different base models (Instruct version) federated fine-tuned on the General NLP challenge, presented in table 3. The errorbar indicates ±$\pm 1$ Std. Dev. on the different downstream tasks.
Figure 9: Accuracy (%) versus system performance for different base models (Instruct version) federated fine-tuned on the Finance challenge, presented in table 4. The errorbar indicates ±$\pm 1$ Std. Dev. on the different downstream tasks.
Figure 10: Accuracy (%) versus system performance for different base models (Instruct version) federated fine-tuned on the Medical challenge, presented in table 5. The errorbar indicates ±$\pm 1$ Std. Dev. on the different downstream tasks.
Figure 11: Pass@1 scores (%) versus system performance for different base models (Instruct version) federated fine-tuned on the Coding challenge, presented in table 6. The errorbar indicates ±$\pm 1$ Std. Dev. on the different downstream tasks.
B.3 Analysis on 24B base models
@@
Models STEM SS Humanities Avg Comm. Mem. PMQA MMCQA MQA CQA Avg Comm. Mem.
@@
Additionally, we extend our investigation to the upper bounds of federated fine-tuning by experimenting with state-of-the-art 24B-parameter models—Mistral-Small-24B-Base-2501 [3] and Mistral-Small-24B-Instruct-2501 [4]—which have been reported to achieve performance comparable to Llama-3.3-70B [4].
B.4 Analysis on domain-specific pre-trained base models

### v2

@@ 6 Conclusion
Limitations and future work. This study focuses primarily on cross-silo federated settings; future work may explore a broader range of scenarios, including more resource-constrained environments and more heterogeneous data distributions. Additionally, the experiments were conducted using a unified configuration following a systematic protocol but without extensive hyper-parameter tuning for each individual model-dataset pair. Performance could likely be improved through more careful hyper-parameter optimization. Furthermore, the experimental results indicate that several classical federated aggregation algorithms (e.g., FedAvg, FedProx) exhibit only marginal performance differences in the context of LLM adapter tuning. This finding highlights the need for the community to develop aggregation algorithms specifically tailored to this parameter space (i.e., low-rank adapters) to more effectively integrate knowledge from diverse clients, representing a promising direction for future research. We envision this work as a catalyst for developing more inclusive, secure, and domain-adapted language models, and invite the broader research community to contribute to and extend this benchmarking effort.
@@ Appendix A Detailed experimental settings
This section provides details on client-level data distribution (Section A.1), the base models (Section A.2), and hyper-parameter configurations (Section A.3) used in the experiments.
A.1 Client-level data distribution
From Table 9, we observe that the number of samples is relatively balanced across clients within each domain – a realistic assumption for cross-institutional FL settings. Variation in instruction and response lengths arises primarily from the inherent properties of the source datasets. We did not partition client data based on class distributions, as the instruction-tuning datasets used in this study are unlabeled. However, the semantic similarity metrics reveal meaningful differences across clients in each domain, particularly in the general NLP, medical, and code domains, where similarity scores approach 0, indicating a notable degree of heterogeneity. The higher similarity values observed in the finance domain are likely inherent to the original dataset prior to splitting.
Table 9: Statistical summary of client-level data distribution for four challenges. “# of Samples” indicates the number of data points per client. “Instruct Avg Length” and “Response Avg Length” represent the average sequence lengths of instructions and responses, respectively, for each client. For “Semantic Similarity,” we compute embeddings for each instruction–response pair using the pre-trained all-MiniLM-L6-v2 model and then average these embeddings for each client. After applying dimensionality reduction, we calculate the cosine similarity between each client and all other clients, reporting the average similarity value per client to illustrate semantic variation (with values closer to 1 indicating greater similarity).
General NLP
# of Samples [2601, 2601, 2600, 2600, 2600, 2600, 2600, 2600, 2600, 2600, 2600, 2600, 2600, 2600, 2600, 2600, 2600, 2600, 2600, 2600]
Instruct Avg Length [59, 59, 59, 60, 59, 59, 59, 59, 60, 59, 59, 60, 59, 59, 60, 58, 59, 60, 59, 59]
Response Avg Length [668, 668, 663, 671, 680, 673, 676, 664, 686, 672, 688, 708, 646, 692, 688, 680, 689, 693, 671, 660]
Semantic Similarity [-0.02, 0.04, 0.03, -0.13, -0.07, 0.03, 0.04, -0.14, -0.0, -0.0, -0.1, -0.14, 0.04, -0.05, -0.04, -0.14, -0.08, -0.14, -0.05, 0.04]
Finance
# of Samples [1536, 1536, 1536, 1536, 1536, 1536, 1536, 1536, 1536, 1536, 1536, 1536, 1536, 1536, 1536, 1536, 1536,
1536, 1536, 1536, 1536, 1536, 1535, 1535, 1535, 1535, 1535, 1535, 1535, 1535, 1535, 1535, 1535, 1535,
1535, 1535, 1535, 1535, 1535, 1535, 1535, 1535, 1535, 1535, 1535, 1535, 1535, 1535, 1535, 1535]
Instruct Avg Length [224, 225, 223, 223, 224, 223, 225, 223, 220, 226, 226, 224, 223, 226, 228, 225, 223, 226, 223, 227, 225, 225, 228, 225, 221,
225, 223, 225, 223, 221, 224, 224, 225, 226, 223, 226, 224, 226, 221, 226, 227, 223, 223, 220, 228, 229, 224, 227, 226, 224]
Response Avg Length [9, 9, 9, 9, 9, 9, 9, 9, 9, 9, 9, 9, 9, 9, 9, 9, 9, 9, 9, 9, 9, 9, 9, 9, 9,
9, 9, 9, 9, 9, 9, 9, 9, 9, 9, 9, 9, 9, 9, 9, 9, 9, 9, 9, 9, 9, 9, 9, 9, 9]
Semantic Similarity [0.72, 0.47, 0.29, 0.65, 0.51, 0.79, 0.79, 0.35, 0.33, 0.67, 0.78, 0.14, 0.79, 0.78, 0.76, 0.7, 0.7,
0.77, 0.71, 0.13, 0.78, 0.58, 0.72, 0.79, 0.64, 0.79, 0.65, 0.63, 0.46, 0.31, 0.79, 0.61, 0.69, 0.79,
0.79, 0.79, 0.77, 0.79, 0.24, 0.41, 0.77, 0.34, 0.77, 0.26, 0.74, 0.76, 0.79, 0.58, 0.77, 0.77]
Medical
# of Samples [1698, 1698, 1698, 1698, 1698, 1698, 1698, 1698, 1698, 1698, 1698, 1698, 1698, 1698, 1698, 1697, 1697, 1697, 1697, 1697]
Instruct Avg Length [92, 92, 92, 90, 92, 93, 94, 91, 92, 91, 92, 91, 91, 92, 92, 92, 91, 93, 92, 92]
Response Avg Length [346, 361, 346, 331, 340, 357, 362, 343, 343, 343, 353, 354, 355, 345, 351, 358, 341, 346, 344, 352]
Semantic Similarity [-0.08, -0.04, -0.02, -0.08, -0.05, -0.03, -0.03, -0.03, -0.05, -0.07, -0.04, -0.07, -0.08, -0.03, -0.02, -0.06, -0.06, -0.09, -0.02, -0.08]
Code
# of Samples [2003, 2003, 2002, 2002, 2002, 2002, 2002, 2002, 2002, 2002]
Instruct Avg Length [96, 96, 101, 97, 97, 99, 97, 98, 98, 98]
Response Avg Length [199, 191, 193, 202, 193, 201, 200, 190, 196, 200]
Semantic Similarity [0.03, -0.1, -0.24, -0.27, -0.05, -0.15, 0.05, 0.03, 0.04, -0.14]
A.2 Base model configurations
@@
Non-Instruct Mistral-7B-v0.3 [mistral7b] 7 B 33
Mistral-Small-24B-Base-2501 [mistral24bbase] 24 B 131
Gemma-2-9B [gemma_2024] 9 B 256
Phi-4 [abdin2024phi] 14 B 100
Llama-3.2-1B [grattafiori2024llama] 1 B 128
Llama-3.2-3B [grattafiori2024llama] 3 B 128
Llama-3.1-8B [grattafiori2024llama] 8 B 128
Qwen2-0.5B [chu2024qwen2] 0.5 B 152
Qwen2.5-1.5B [yang2024qwen2] 1.5 B 152
Qwen2.5-3B [yang2024qwen2] 3 B 152
Qwen2.5-7B [yang2024qwen2] 7 B 152
SmolLM2-135M [allal2025smollm2] 135 M 49
SmolLM2-360M [allal2025smollm2] 360 M 49
SmolLM2-1.7B [allal2025smollm2] 1.7 B 49
Instruct Mistral-7B-Instruct-v0.3 [mistral7binstruct] 7 B 33
Mistral-Small-24B-Instruct-2501 [mistral24binstruct] 24 B 131
Gemma-2-9B-Instruct [gotocompany2024gemma2] 9 B 256
Phi-4-Mini-Instruct [abouelenin2025phi] 3.8 B 200
Llama-3.2-1B-Instruct [grattafiori2024llama] 1 B 128
Llama-3.2-3B-Instruct [grattafiori2024llama] 3 B 128
Llama-3.1-8B-Instruct [grattafiori2024llama] 8 B 128
Qwen2-0.5B-Instruct [chu2024qwen2] 0.5 B 152
Qwen2.5-1.5B-Instruct [yang2024qwen2] 1.5 B 152
Qwen2.5-3B-Instruct [yang2024qwen2] 3 B 152
Qwen2.5-7B-Instruct [yang2024qwen2] 7 B 152
SmolLM2-135M-Instruct [allal2025smollm2] 135 M 49
SmolLM2-360M-Instruct [allal2025smollm2] 360 M 49
SmolLM2-1.7B-Instruct [allal2025smollm2] 1.7 B 49
A.3 Hyper-parameter configurations
@@
Appendix B Additional experimental results
B.1 Federated fine-tuning on non-instruct pre-trained base models
@@
Models STEM $S$ Humanities Avg Comm. Mem. PMQA MMCQA MQA CQA Avg Comm. Mem.
@@
B.2 Pareto analysis of evaluation performance
Figure 4: Accuracy (%) versus system performance for different non-instruct base models federated fine-tuned on the General NLP challenge, presented in table 12. The errorbar indicates $\pm 1$ Std. Dev. on the different downstream tasks.
Figure 5: Accuracy (%) versus system performance for different non-instruct base models federated fine-tuned on the Finance challenge, presented in table 13. The errorbar indicates $\pm 1$ Std. Dev. on the different downstream tasks.
Figure 6: Accuracy (%) versus system performance for different non-instruct base models federated fine-tuned on the Medical challenge, presented in table 14. The errorbar indicates $\pm 1$ Std. Dev. on the different downstream tasks.
Figure 7: Pass@1 scores (%) versus system performance for different non-instruct base models federated fine-tuned on the Coding challenge, presented in table 15. The errorbar indicates $\pm 1$ Std. Dev. on the different downstream tasks.
Figure 8: Accuracy (%) versus system performance for different base models (Instruct version) federated fine-tuned on the General NLP challenge, presented in table 3. The errorbar indicates $\pm 1$ Std. Dev. on the different downstream tasks.
Figure 9: Accuracy (%) versus system performance for different base models (Instruct version) federated fine-tuned on the Finance challenge, presented in table 4. The errorbar indicates $\pm 1$ Std. Dev. on the different downstream tasks.
Figure 10: Accuracy (%) versus system performance for different base models (Instruct version) federated fine-tuned on the Medical challenge, presented in table 5. The errorbar indicates $\pm 1$ Std. Dev. on the different downstream tasks.
Figure 11: Pass@1 scores (%) versus system performance for different base models (Instruct version) federated fine-tuned on the Coding challenge, presented in table 6. The errorbar indicates $\pm 1$ Std. Dev. on the different downstream tasks.
B.3 Analysis on 24B base models
Additionally, we extend our investigation to the upper bounds of federated fine-tuning by experimenting with state-of-the-art 24B-parameter models—Mistral-Small-24B-Base-2501 [mistral24bbase] and Mistral-Small-24B-Instruct-2501 [mistral24binstruct]—which have been reported to achieve performance comparable to Llama-3.3-70B [mistral24binstruct].
B.4 Analysis on domain-specific pre-trained base models